OpenStudy (anonymous):

pl help asap gold medal will be awarded.generate the first 5 terms of this sequence. f(1)=1and f(2)=2, f(n-1) + f(n-2) , for n>2

OpenStudy (anonymous):

can u help? plz

OpenStudy (plainntall):

so we now that f(1)=1 and f(2)=2 for n>2 we have n=3 f(3-1)+(f(3-2)=f(2)+f(1)=?

OpenStudy (anonymous):

um i dont know

OpenStudy (plainntall):

f(2)+f(1) = what? remember we know that f(1)=1 and f(2)=2

OpenStudy (anonymous):

um 2+1?

OpenStudy (plainntall):

yes so now we know that f(3)=3 so what happens for n=4? f(n-1)+f(n-2)=?

OpenStudy (anonymous):

we plug it in

OpenStudy (plainntall):

right f(4-1)+f(4-2)=f(3)+f(2)=?

OpenStudy (plainntall):

We just figured out that f(3)=3 and we were give that f(2)=2

OpenStudy (anonymous):

f(3)+F(2+F(3)

OpenStudy (plainntall):

f(4)=f(3)+f(2)=3+2=?

OpenStudy (anonymous):

do add just the back part

OpenStudy (plainntall):

You figure out what the function is for each value and put it into the equation.

OpenStudy (anonymous):

um am so sorry to interupt but am really in a hurry so can you plz give me final answer

OpenStudy (plainntall):

we know that the first terms are (1,2,3,5) for terms 1,2,3,4 each new term is equal to the sum of the function at n-1 and n-2 so for n=5 substitute n into f(n-1)+f(n-2) what do you get? f(?) + f(?)

OpenStudy (anonymous):

i dont kow how to solve it

OpenStudy (plainntall):

If you substitute n=5 you get f(5-1) +f(5-2) which is f(4)+f(3) we evaluated the function at 3 and 4 already so we have f(4)+f(3)= 5 +3 = 8 so (1,2,3,5,8) are the first 5 terms

OpenStudy (anonymous):

oh ok i get it

OpenStudy (anonymous):

thx